This last week the two engineering teams from Austin worked both days and nights to get their work done. The day time team was able to get the mounts for the Tip-Tilt camera, VAT and DMI on the telescope. They were able to get the VAT installed and found that the center mirror is just 30 mm off from the center of the X and Y tracker encoder ranges. We still need to determine where the encoder ranges are with respect to the upper hex which brings us to the work of the other engineering team. This team worked most of the nights this past week including weekends to get the 97 point test started. They determined that to take a set of 18 laser tracker data points at each tracker position takes far longer than anticipated and so descoped to a 21 point test instead of a 91 point test. Even with this descope the 21 point test took 15 hours of work.
This week the Austin engineering teams have departed and are working on new mounts for the test mass based on lessons learned from this past week. The data from the 21 point test needs to be analyzed and we will followup any questions that analysis may bring up with further laser tracker measurements.
The mirror teams are swapping two mirrors into the array this week and later this week the cabling teams will be putting on the clamps in the Azimuth cable wrap and fixing any cabling problems that might have developed with the full range to tracker motion that we can now make.
The contractors from Veliz Construction made substantial progress this past week. They completed the first concrete pour for the pipe and electrical vault and this week they will be setting up the forms for the second pour (which will be the walls of the vault. We will also have the electrical contractor on site to complete some Phase I punch list items and review the site for the Phase II electrical work. We will also have Gaffey Crane on site to work on the dome crane box which must have a new form factor to avoid collisions with the stinger from the Y-drive.